OpenStudy (anonymous):

Solve the system: -x-y-3z=-9 z=-3x-1 x=5y-z+23

OpenStudy (anonymous):

(But don't actually solve it for me, just help me out?)

OpenStudy (usukidoll):

well you can either plug in z from the second equation into the first equation...

OpenStudy (08surya):

put the value of z in last equation then it it will convert into x and y then put value of z in first equation it will also get converted into two variable x and y. now u have two equations with two unknowns .solve it

OpenStudy (usukidoll):

or grab the x equation and subsitute it into the first equation.

OpenStudy (anonymous):

i did that andi ended up with x=8y+24 and y+8x=-12 and i tried to combine them and eliminate 8x but that didnt go well

OpenStudy (08surya):

multiply with 8 in first equation that u got and multiply 1 with second equation now substract both equation u will get value of y

OpenStudy (anonymous):

(x=8y+24)8 y+8x=-12 8x=64y+192 -8x=y+12 I'd end up with 65y+204

OpenStudy (08surya):

no u have end up with 65y+204=0 so 65y=-204 y=-204/65

OpenStudy (anonymous):

i'm not supposed to end up with fractions

OpenStudy (08surya):

so what r u getting

OpenStudy (anonymous):

well i'm gunna try again another way

OpenStudy (08surya):

ur choice